# Computation of Smyth and Hoare Power Constructions in Well-filtered Dcpos

### v2

@@ 3.2 The Smyth power construction
Xu et al. proved that for a well-filtered space $X$, $\Sigma\mathcal{Q}(X)$ is still well-filtered [24], which indicates that for each well-filtered dcpo $L$, $\mathcal{Q}(L)$ with the Scott topology is well-filtered, that is, $\mathcal{Q}(L)$ is an object in $\mathbf{WF}$. This fact makes us give a rational Smyth power construction.
@@
The Smyth power construction for a well-filtered dcpo is the set of all nonempty Scott compact saturated subsets ordered by the reverse inclusion.
We shall use a direct way being different from that of $\mathcal{H}$ to establish $\mathcal{Q}$ as a monad on $\mathbf{WF}$.
@@
Let $f:L\rightarrow M$ be a Scott-continuous mapping between well-filtered dcpos $L$ and $M$. Then $\mathcal{Q}(f):\mathcal{Q}(L)\rightarrow\mathcal{Q}(M)$ defined as
@@
Proof.
Let $\{K_{i}:i\in I\}$ be a filtered family in $\mathcal{Q}(L)$ with $\sup_{i\in I}K_{i}$ existing, that is to say, $\bigcap_{i\in I}K_{i}\in\mathcal{Q}(L)$. Since $f$ is Scott-continuous, $\{{\uparrow}f(K_{i}):i\in I\}$ is a subset of $\mathcal{Q}(M)$ and filtered. By the well-filteredness of $M$, we know $\bigcap_{i\in I}{\uparrow}f(K_{i})\in\mathcal{Q}(M)$, which means $\sup_{i\in I}{\uparrow}f(K_{i})$ exists in $\mathcal{Q}(M)$. Now we just need to prove that $\mathcal{Q}(f)(\bigcap_{i\in I}K_{i})=\bigcap_{i\in I}{\uparrow}f(K_{i})$, more precisely, to show
$\bigcap_{i\in I}{\uparrow}f(K_{i})\subseteq\mathcal{Q}(f)(\bigcap_{i\in I}K_{i})={\uparrow}f(\bigcap_{i\in I}K_{i})$
holds since the other containment is clear. If there is an element $x\in\bigcap_{i\in I}{\uparrow}f(K_{i})$ but $x\notin{\uparrow}f(\bigcap_{i\in I}K_{i})$, then $\bigcap_{i\in I}K_{i}$ will be contained in $f^{-1}(M\setminus{\downarrow}x)\in\sigma(L)$. Due to the well-filteredness of $L$, there must be an $i_{0}\in I$ such that $K_{i_{0}}\subseteq f^{-1}(M\setminus{\downarrow}x)$. Thus $f(K_{i_{0}})\subseteq M\setminus{\downarrow}x$ and so $x\notin{\uparrow}f(K_{i_{0}})$, which contradicts the assumption that $x\in\bigcap_{i\in I}{\uparrow}f(K_{i})$. ∎
@@
The saturation is clear. We only consider the compactness. Let $\mathcal{U}$ be a directed subset of $\sigma(L)$ with $\bigcup\mathcal{K}\subseteq\bigcup\mathcal{U}$. Then for each $K\in\mathcal{K}$, its compactness yields the existence of $U_{K}\in\mathcal{U}$ such that $K\subseteq U_{K}$, that is, $K\in\Box U_{K}$. Since the well-filteredness of $L$ hints that each $\Box U$ is Scott open in $\mathcal{Q}(L)$ and $\mathcal{U}$ is directed, $\{\Box U:U\in\mathcal{U}\}$ forms a directed family of Scott open subsets of $\mathcal{Q}(L)$ containing $\mathcal{K}$. So $\mathcal{K}\subseteq\Box U_{0}$ for some $U_{0}\in\mathcal{U}$ by its compactness. Thus $\bigcup\mathcal{K}\subseteq U_{0}$. ∎
@@
The endofunctor $\mathcal{Q}$ with the unit $\theta$ and the multiplication $\iota$ forms a monad on $\mathbf{WF}$, where $\theta$ and $\iota$ are defined concretely as follows when they work on a well-filtered dcpo $L$,
@@
$\iota_{L}(\mathcal{K})=\bigcup\mathcal{K}$, for any $\mathcal{K}\in\mathcal{Q}((\mathcal{Q}(L)))$.
@@
It is a routine to check that $\mathcal{Q}$ is a monad. ∎
@@
The definitions of $\mathcal{Q}$-algebras reveal that $M$ is a retract of $\mathcal{Q}(M)$. Then by [20, proposition 5.6], we know that $M$ is a deflationary semilattice in $\mathbf{WF}$.
@@
Proof.
It is easy to prove that $\alpha(K)=\wedge K$ by utilizing the condition $\alpha\circ\theta_{M}=id_{M}$. ∎
@@
Each $\mathcal{Q}$-algebra homomorphism preserves finite infs.
Proof.
Let $h:(L,\alpha_{L})\rightarrow(M,\alpha_{M})$ be a $\mathcal{Q}$-algebra homomorphism. Then we have $h\circ\alpha_{L}(K)=\alpha_{M}\circ\mathcal{Q}(h)(K)$ for any $K\in\mathcal{Q}(L)$, that is, $h(\wedge K)=\wedge h(K)$ holds. Accordingly, one can deduce that $h$ preserves finite infs as every finitely generated subset is compact saturated. ∎
@@
Unfortunately, we can not accurately characterize what the $\mathcal{Q}$-algebras are. What is more, $\mathcal{Q}$ can not even give the free deflationary semilattice for $\mathbf{WF}$. Once $\mathcal{Q}$ can, each free deflationary semilattice in $\mathbf{WF}$ together with a structure map will be a $\mathcal{Q}$-algebra, which is violated in the following counterexample.
@@
=$\mathcal{L}=\{A\subseteq[0,1]:A\neq\emptyset\ \mathrm{and}\ [0,1]\setminus A\ \mathrm{is\ countable}\}$.
It has been proved to be a deflationary semilattice in $\mathbf{DCPO}$. Besides, we can see that every nonempty finite subset of $\mathcal{L}$ has a sup, which is indeed the union of its elements. It then follows that $\mathcal{L}$ is an up-complete $\vee$-semilattice, this implies the well-filteredness of $\Sigma\mathcal{L}$ witnessed by [23, Proposition 3.1]. Thus $\mathcal{L}$ is a deflationary semilattice in $\mathbf{WF}$. However, there exists a compact saturated subset
$K=\{[0,1]\setminus\{x\}:x\in[0,1/2]\}$,
@@ 4 The commutativity between $\mathcal{H}$ and $\mathcal{Q}$
For convenience, we will omit the subscripts when the object acted on is clear.
@@
In [23], Corollary 3.2 guarantees the well-filteredness of a complete lattice $L$. Since for any $x,y\in L$, ${\uparrow}x\cap{\uparrow}y={\uparrow}(x\vee y)$, by Corollary 3.2 in [15], we have that $L$ is coherent. ∎
@@
The functions $\phi$ and $\psi$ are well-defined.
@@
for any $K\in\mathcal{Q}(L)$, $\phi^{\prime}(K)=\{A\in\mathcal{H}(L):A\cap K\neq\emptyset\}$,
for any $A\in\mathcal{H}(L)$, $\psi^{\prime}(A)=\{K\in\mathcal{Q}(L):K\cap A\neq\emptyset\}$.
Then $\phi(\mathcal{A})=\bigcap\{\phi^{\prime}(K):K\in\mathcal{A}\}$ and $\psi(\mathcal{K})=\bigcap\{\psi^{\prime}(A):A\in\mathcal{K}\}$. One can easily check that $\phi^{\prime}(K)={\uparrow}\eta_{L}(K)$ and $\psi^{\prime}(A)=cl(\theta_{L}(A))$, which implies that $\phi^{\prime}$ and $\psi^{\prime}$ are well-defined. We know the intersection of arbitrary closed subsets is still closed. So, $\psi$ is well-defined. Now we focus on $\phi$. Lemma 4.2 reveals that $\Sigma\mathcal{H}(L)$ is coherent. Combining with the well-filteredness of $\Sigma\mathcal{H}(L)$, it will be a fact that each intersection of nonempty collection of compact saturated subsets in $\mathcal{H}(L)$ is still compact saturated. Thus $\phi(\mathcal{A})\in\mathcal{Q}(\mathcal{H}(L))$, i.e., $\phi$ is well-defined.
@@
A poset $P$ ($T_{0}$ space $X$) is said to have property $(\mathrm{KC})$ if for every compact saturated subset $\mathcal{K}$ of $\Gamma(P)$ under the Scott topology ($\mathcal{K}$ of $\Gamma(X)$ with respect to the lower Vietoris topology), any $U\in\sigma(P)$ ($U\in\mathcal{O}(X)$) that meets all members of $\mathcal{K}$ contains a compact saturated subset $K$ that still meets all members of $\mathcal{K}$.
Recall that a space $X$ is consonant if and only if for any Scott open subset $\mathcal{U}$ in $\mathcal{O}(X)$ and any $U\in\mathcal{U}$, there exists a $K\in Q(X)$ such that $U\in\Phi(K)\subseteq\mathcal{U}$, where $\Phi(K)=\{V\in\mathcal{O}(X):K\subseteq V\}$. The concept was proposed initially by Dolecki et al. in [5] to answer some questions related to topologies on the hyperspace of closed subsets of a topological space. We say a poset consonant if it endowed with the Scott topology is consonant.
@@
Let $\mathcal{K}$ be a compact saturated subset of $(\Gamma(P),\sigma(\Gamma(P)))$ and $U$ a Scott open subset of $P$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Set $\mathcal{U}=\bigcap\{\vartriangle\!\!A:A\in\mathcal{K}\}$, where $\vartriangle\!\!A=\{V\in\sigma(P):V\cap A\neq\emptyset\}$. We claim that $\mathcal{U}\in\sigma(\sigma(P))$. Obviously, it is an upper set. Assume $\mathcal{V}$ is a directed subset of $\sigma(P)$ and $\sup\mathcal{V}=\bigcup\mathcal{V}\in\mathcal{U}$. Then $\bigcup\mathcal{V}\in\vartriangle\!\!A$ for all $A\in\mathcal{K}$, that is, $\bigcup\mathcal{V}\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. It means that for each $A\in\mathcal{K}$, there exists $V_{A}\in\mathcal{V}$ such that $V_{A}\cap A\neq\emptyset$, i.e., $A\in\lozenge V_{A}$, and thus $\mathcal{K}\subseteq\bigcup\{\lozenge V:V\in\mathcal{V}\}$. By compactness of $\mathcal{K}$ and directedness of $\{\lozenge V:V\in\mathcal{V}\}$, there must exist a $V_{0}\in\mathcal{V}$ such that $\mathcal{K}\subseteq\lozenge V_{0}$. It follows that $V_{0}\in\vartriangle\!\!A$ for each $A\in\mathcal{K}$, that is to say, $V_{0}\in\mathcal{U}$. Thus $\mathcal{U}$ is Scott open. Since $P$ is consonant and $U\in\mathcal{U}$, there exists a $K\in Q(P)$ satisfying $U\in\Phi(K)\subseteq\mathcal{U}$. We will complete the proof if we show that $K\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Suppose $K\cap A_{0}=\emptyset$ for some $A_{0}\in\mathcal{K}$. Then $K\subseteq P\setminus A_{0}$, which implies $P\setminus A_{0}\in\Phi(K)\subseteq\mathcal{U}$; so $P\setminus A_{0}\in\vartriangle\!\!A_{0}$, a contradiction. ∎
Lemma 4.6.
A $T_{0}$ space $X$ is consonant if and only if it has the property $(\mathrm{KC})$.
Proof.
()($\Rightarrow$)( ⇒ ): Let $\mathcal{K}$ be a compact saturated subset of $(\Gamma(X),\nu(\Gamma(X)))$ and $U\in\mathcal{O}(X)$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Set $\mathcal{U}=\bigcap\{\blacktriangle A:A\in\mathcal{K}\}$, where $\blacktriangle A=\{V\in\mathcal{O}(X):V\cap A\neq\emptyset\}$. Then similar to the proof of Lemma 4.5, one can easily verify that $\mathcal{U}\in\sigma(\mathcal{O}(X))$ and find some $K\in Q(X)$ qualified.
()($\Leftarrow$)( ⇐ ): Let $\mathcal{V}$ be a Scott open subset of $\mathcal{O}(X)$ and $V\in\mathcal{V}$. We set $\mathcal{K}=\bigcap\{\lozenge U:U\in\mathcal{V}\}$. Notably, $\mathcal{K}$ is indeed the image of $\mathcal{V}$ under the function $\psi_{X}:\sigma(\mathcal{O}(X))\rightarrow\mathcal{Q}(\Gamma(X))$ given in [4, Definition 6.7], where $\mathcal{Q}(\Gamma(X))$ is the set of all compact saturated subset of $\Gamma(X)$ with respect to the lower Vietoris topology. Thus $\mathcal{K}$ is a compact saturated subset of $(\Gamma(X),\nu(\Gamma(X)))$. As $V\in\mathcal{V}$ suggests that $\mathcal{K}\subseteq\lozenge V$, $X\setminus V\notin\mathcal{K}$, more precisely, each $A$ in $\mathcal{K}$ is not contained in $X\setminus V$, that is, $V\cap A\neq\emptyset$ for each $A\in\mathcal{K}$. Since $X$ has the property $(\mathrm{KC})$, we can find a $K\in Q(X)$ contained in $V$ that intersects each $A\in\mathcal{K}$. To make the proof complete, we only need to show that $\Phi(K)\subseteq\mathcal{V}$. For each $W\in\Phi(K)$, the fact $K\subseteq W$ yields that $W\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. So $\mathcal{K}\in\Box\lozenge W$, which is equivalent to saying $W\in\phi_{X}({\mathcal{K}})$, where $\phi_{X}:\mathrm{K}(\Gamma(X))\rightarrow\sigma(\mathcal{O}(X))$ is also given in [4, Definition 6.7]. Moreover, [4, Theorem 6.10] has shown that $\phi_{X}\circ\psi_{X}=id_{\sigma(\mathcal{O}(X))}$, which enables us to deduce that $W\in\phi_{X}({\mathcal{K}})=\phi_{X}(\psi_{X}(\mathcal{V}))=\mathcal{V}$. Hence, we have $V\in\Phi(K)\subseteq\mathcal{V}$, which indicates the consonance of $X$.
Now see [4, Theorem 6.10] again, it states that for a space $X$, $\sigma(\mathcal{O}(X))$ and $\mathcal{Q}(\Gamma(X))$ are isomorphic lattices. Recently, Chen et al. in [3] proved that if a poset $P$ under the Scott topology is core-compact, then $\sigma(\Gamma(P))=\nu(\Gamma(P))$. Thus for a poset $P$, $\sigma(\sigma(P))$ would be isomorphic to $Q(\Gamma(P))$ as long as $\Sigma P$ is core-compact. Combining all of these results and the above two lemmas, we obtain the following result.
@@
()($\Rightarrow$)( ⇒ ): Let $\mathcal{U}$ be a Scott open subset of $\mathcal{Q}(L)$ and $K$ an element in $\mathcal{U}$. Then $K$ does not belong to the Scott closed subset $\mathcal{Q}(L)\setminus\mathcal{U}$. Since $\psi$ is surjective, there exists a $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$ such that $\mathcal{Q}(L)\setminus\mathcal{U}=\psi(\mathcal{K})$, then $K\notin\psi(\mathcal{K})$. By the definition of $\psi$, it yields that $K\cap A=\emptyset$ for some $A\in\mathcal{K}$. Now we find a Scott open subset $V=L\setminus A$ containing $K$, that is, $K\in\Box V$. Thus we only need to show $\Box V\subseteq\mathcal{U}$. For any $M\in\Box V$, $M\cap A=\emptyset$, which implies that $M\notin\psi(\mathcal{K})$. Hence, $M\in\mathcal{U}$.
Let $\mathcal{K}$ be a Scott compact saturated subset of $\Gamma(L)$ and $U\in\sigma(L)$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Then $L\setminus U\notin\mathcal{K}$; otherwise, by the saturation of $\mathcal{K}$, there must be an $A_{0}\in\mathcal{K}$ such that $A_{0}\subseteq L\setminus U$, a contradiction. The fact that $\phi$ is surjective suggests that there is an $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ such that $\mathcal{K}=\phi(\mathcal{A})$. Thus $L\setminus U\notin\phi(\mathcal{A})$. This means $(L\setminus U)\cap K_{0}=\emptyset$ i.e., $K_{0}\subseteq U$ for some $K_{0}\in\mathcal{A}$. Now suppose that $K_{0}\cap A_{1}=\emptyset$ for some $A_{1}\in\mathcal{K}$. Then $A_{1}\notin\phi(\mathcal{A})=\mathcal{K}$, which is an contradiction obviously. Thus $K_{0}$ intersects each $A\in\mathcal{K}$ and $L$ has the property $(\mathrm{KC})$ as desired.
()($\Leftarrow$)( ⇐ ): Obviously, $\phi$ and $\psi$ are order-preserving. Thus this proof will be completed so long as we can check that $\psi\circ\phi$ and $\phi\circ\psi$ are identities, more precisely, the following two equations hold (for arbitrary $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ and $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$):

### v3

@@ 3.2 The Smyth power construction
Xu et al. proved that for a well-filtered space $X$, $\Sigma\mathcal{Q}(X)$ is still well-filtered [26, Theorem 6.5], which indicates that for each well-filtered dcpo $L$, $\mathcal{Q}(L)$ with the Scott topology is well-filtered, that is, $\mathcal{Q}(L)$ is an object in $\mathbf{WF}$. This fact makes us give a well-defined Smyth power construction.
@@
The Smyth power construction for a well-filtered dcpo is the set of all Scott compact saturated subsets ordered by reverse inclusion.
Recall that, inspired by the facts that for a sober space $X$, the set $\mathcal{Q}(X)$ ordered by reverse inclusion is isomorphic to the set of Scott open filters in $\mathcal{O}(X)$ ordered by inclusion, and for a dcpo $L$, the set of all compact saturated subsets $\mathcal{Q}(L)$ ordered by reverse inclusion is no longer a dcpo, Andrea Schalk focused on the set of all Scott open filters $OFilt(\sigma(L))$, proved that $OFilt(\sigma(L))$ ordered by inclusion remains a dcpo. Concerning this, she gave an endofunctor on the category $\mathbf{DCPO}$ and yields a monad (one can refer to [21, Lemma 7.5, Theorem 7.7]). Fortunately, for a well-filtered dcpo $L$, the set of all Scott compact saturated subsets ordered by reverse inclusion remains a well-filtered dcpo as shown in [26]. Next we will construct an endofunctor $\mathcal{Q}$ on $\mathbf{WF}$ and verify that it is also a monad.
@@
[11, Theorem 3.9] Let $f:L\rightarrow M$ be a Scott-continuous mapping between well-filtered dcpos $L$ and $M$. Then $\mathcal{Q}(f):\mathcal{Q}(L)\rightarrow\mathcal{Q}(M)$ defined as
@@
The saturation is clear. We only consider the compactness. Let $\mathcal{U}$ be a directed subset of $\sigma(L)$ with $\bigcup\mathcal{K}\subseteq\bigcup\mathcal{U}$. Then for each $K\in\mathcal{K}$, its compactness yields the existence of $U_{K}\in\mathcal{U}$ such that $K\subseteq U_{K}$, that is, $K\in\Box U_{K}$. Since the well-filteredness of $L$ implies that each $\Box U$ is Scott open in $\mathcal{Q}(L)$ and $\mathcal{U}$ is directed, $\{\Box U:U\in\mathcal{U}\}$ forms a directed family of Scott open subsets of $\mathcal{Q}(L)$ containing $\mathcal{K}$. So $\mathcal{K}\subseteq\Box U_{0}$ for some $U_{0}\in\mathcal{U}$ by its compactness. Thus $\bigcup\mathcal{K}\subseteq U_{0}$.
@@
The endofunctor $\mathcal{Q}$ with the unit $\theta$ and the multiplication $\iota$ forms a monad on $\mathbf{WF}$, where $\theta$ and $\iota$ are defined as follows when they work on a well-filtered dcpo $L$,
@@
$\iota_{L}(\mathcal{K})=\bigcup\mathcal{K}$, for any $\mathcal{K}\in\mathcal{Q}(\mathcal{Q}(L))$.
@@
It is routine to check that $\mathcal{Q}$ is a monad.
@@
The definitions of $\mathcal{Q}$-algebras reveal that $M$ is a retract of $\mathcal{Q}(M)$. Then by [21, proposition 5.6], we know that $M$ is a deflationary semilattice in $\mathbf{WF}$.
Based on [21, Lemma 4.4] given by Andrea Schalk, we can derive the following results.
@@
Each $\mathcal{Q}$-algebra homomorphism preserves infs of compact sets, hence in particular finite infs.
@@
Unfortunately, we cannot accurately characterize what the $\mathcal{Q}$-algebras are. Moreover, $\mathcal{Q}$ does not give the free deflationary semilattice for $\mathbf{WF}$ because there are deflationary semilattices in $\mathbf{WF}$ that are not $\mathcal{Q}$-algebras, as shown in the following example.
@@
$\mathcal{L}=\{A\subseteq[0,1]:A\neq\emptyset\ \mathrm{and}\ [0,1]\setminus A\ \mathrm{is\ countable}\}$.
It has been proved to be a deflationary semilattice in $\mathbf{DCPO}$. Besides, we can see that every nonempty finite subset of $\mathcal{L}$ has a sup, which is indeed the union of its elements. It then follows that $\mathcal{L}$ is an up-complete $\vee$-semilattice, this implies the well-filteredness of $\Sigma\mathcal{L}$ witnessed by [25, Proposition 3.1]. Thus $\mathcal{L}$ is a deflationary semilattice in $\mathbf{WF}$. However, there exists a compact saturated subset
$K={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\uparrow}}\{[0,1]\setminus\{x\}:x\in[0,1/2]\}$,
@@ 4 The commutativity between $\mathcal{H}$ and $\mathcal{Q}$
Notably, the forms of these two definitions are similar to those in [4] and [11], where the objects they discussed are quasi-polish spaces and $\mathcal{U}_{S}$-admitting dcpos, respectively. For convenience, we will omit the subscripts when the object acted on is clear.
@@
[25, Corollary 3.2] guarantees the well-filteredness of a complete lattice $L$. Since for any $x,y\in L$, ${\uparrow}x\cap{\uparrow}y={\uparrow}(x\vee y)$, by [15, Corollary 3.2], we have that $L$ is coherent.
@@
The functions $\phi$ and $\psi$ defined in Definition 4.1 are well-defined.
@@
$\phi^{\prime}(K)=\{A\in\mathcal{H}(L):A\cap K\neq\emptyset\}$, for any $K\in\mathcal{Q}(L)$,
$\psi^{\prime}(A)=\{K\in\mathcal{Q}(L):K\cap A\neq\emptyset\}$, for any $A\in\mathcal{H}(L)$.
Then $\phi(\mathcal{A})=\bigcap\{\phi^{\prime}(K):K\in\mathcal{A}\}$ and $\psi(\mathcal{K})=\bigcap\{\psi^{\prime}(A):A\in\mathcal{K}\}$. One can easily check that $\phi^{\prime}(K)=\mathcal{Q}(\eta_{L})(K)={\uparrow}\eta_{L}(K)$ and $\psi^{\prime}(A)=\mathcal{H}(\theta_{L})(A)=cl(\theta_{L}(A))$, which implies that $\phi^{\prime}$ and $\psi^{\prime}$ are well-defined. We know the intersection of arbitrary closed subsets is still closed. So, $\psi$ is well-defined. Now we focus on $\phi$. Lemma 4.2 reveals that $\Sigma\mathcal{H}(L)$ is coherent. Combining with the well-filteredness of $\Sigma\mathcal{H}(L)$, it will be a fact that each intersection of a collection of compact saturated subsets in $\mathcal{H}(L)$ is still compact saturated. Thus $\phi(\mathcal{A})\in\mathcal{Q}(\mathcal{H}(L))$, i.e., $\phi$ is well-defined.
Proposition 4.4.
Let $L$ be a well-filtered dcpo. The function $\phi:\mathcal{HQ}(L)\rightarrow\mathcal{QH}(L)$ defined in Definition 4.1 is Scott-continuous.
Proof.
Assume that there are $\mathcal{A},\mathcal{B}\in\mathcal{HQ}(L)$ with $\mathcal{A}\leq\mathcal{B}$, i.e., $\mathcal{A}\subseteq\mathcal{B}$. By the definition of $\phi$, one can easily see that $\phi(\mathcal{B})\subseteq\phi(\mathcal{A})$, equivalently, $\phi(\mathcal{A})\leq\phi(\mathcal{B})$. So, $\phi$ is order preserving. Let $\{\mathcal{A}_{i}:i\in I\}$ be a directed subset of $\mathcal{HQ}(L)$. To complete this proof, we need to verify that $\phi(\bigvee_{i\in I}\mathcal{A}_{i})=\bigvee_{i\in I}\phi(\mathcal{A}_{i})$, i.e., $\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))=\bigcap_{i\in I}\phi(\mathcal{A}_{i})$. Since $\phi$ is order preserving, $\bigcap_{i\in I}\phi(\mathcal{A}_{i})\subseteq\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$ clearly holds. Now pick $A\in\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$, then for each $K\in\bigcup_{i\in I}\mathcal{A}_{i},A\cap K\neq\emptyset$. So we have $A\cap K\neq\emptyset,\forall i\in I,\forall K\in\mathcal{A}_{i}$, which immediately implies that $A\in\phi(\mathcal{A}_{i}),\forall i\in I$. Therefore, $A\in\bigcap_{i\in I}\phi(\mathcal{A}_{i})$ consequently, $\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))\subseteq\bigcup_{i\in I}\phi(\mathcal{A}_{i})$ holds.
@@
A poset $P$ is said to have property $(\mathrm{KC})$ if for every compact saturated subset $\mathcal{K}$ of $\Gamma(P)$ under the Scott topology, any $U\in\sigma(P)$ that meets all members of $\mathcal{K}$ contains a compact saturated subset $K$ that still meets all members of $\mathcal{K}$.
Recall that a space $X$ is consonant if and only if for any Scott open subset $\mathcal{U}\in\mathcal{O}(X)$ and any $U\in\mathcal{U}$, there exists a $K\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)}$ such that $U\in\Phi(K)\subseteq\mathcal{U}$, where $\Phi(K)=\{V\in\mathcal{O}(X):K\subseteq V\}$. The concept was proposed initially by Dolecki et al. in [5] to answer some questions related to topologies on the hyperspace of closed subsets of a topological space. We say a poset is consonant if it endowed with the Scott topology is consonant.
@@
Let $\mathcal{K}$ be a compact saturated subset of $(\Gamma(P),\sigma(\Gamma(P)))$ and $U$ a Scott open subset of $P$ satisfying that $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Then for each $A\in\mathcal{K}$, $A\in\Diamond U$, which implies that $\mathcal{K}\subseteq\Diamond U$, i.e., $\mathcal{K}\in\Box\Diamond U$. Notably, [4, Theorem 6.13] tells us that the consonance of a topological space $X$ implies the equation $\tau^{-1}(\Diamond\Box U)=\Box\Diamond U$ holds. For the space $\Sigma P$, $\tau$ in [4, Theorem 6.13] is just $\psi$ defined by us. So the consonance of $\Sigma P$ will indicate that $\Box\Diamond U\subseteq\psi^{-1}(\Diamond\Box U)$. Thus $\mathcal{K}\in\psi^{-1}(\Diamond\Box U)$. From which we have that $\psi(\mathcal{K})\in\Diamond\Box U$, i.e., $\psi(\mathcal{K})\cap\Box U\neq\emptyset$. So there exists a $K\in\psi(\mathcal{K})$ contained in $U$. Therefore, we find a compact saturated subset $K$ contained in $U$ that meets each element of $\mathcal{K}$. Hence, $P$ has the property $(\mathrm{KC})$.
Now see [4, Theorem 6.10] again, it states that for a space $X$, $\sigma(\mathcal{O}(X))$ and $\mathcal{Q}(\Gamma(X))$ are isomorphic lattices. Recently, Chen et al. in [3] proved that if a poset $P$ under the Scott topology is core-compact, then $\sigma(\Gamma(P))=\nu(\Gamma(P))$. Thus for a poset $P$, $\sigma(\sigma(P))$ would be isomorphic to $\mathcal{Q}(\Gamma(P))$ as long as $\Sigma P$ is core-compact. Combining all of these results and the above lemma, we obtain the following result.
@@
$(\Rightarrow)$: Let $\mathcal{U}$ be a Scott open subset of $\mathcal{Q}(L)$ and $K$ an element in $\mathcal{U}$. Then $K$ does not belong to the Scott closed subset $\mathcal{Q}(L)\setminus\mathcal{U}$. Since $\psi$ is surjective, there exists a $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$ such that $\mathcal{Q}(L)\setminus\mathcal{U}=\psi(\mathcal{K})$, then $K\notin\psi(\mathcal{K})$. By the definition of $\psi$, it yields that $K\cap A=\emptyset$ for some $A\in\mathcal{K}$. The Scott open subset $V=L\setminus A$ contains $K$, that is, $K\in\Box V$. Thus we only need to show $\Box V\subseteq\mathcal{U}$. For any $M\in\Box V$, $M\cap A=\emptyset$, which implies that $M\notin\psi(\mathcal{K})$. Hence, $M\in\mathcal{U}$.
Let $\mathcal{K}$ be a Scott compact saturated subset of $\Gamma(L)$ and $U\in\sigma(L)$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Because of $U\cap(L\setminus U)=\emptyset$, we have $L\setminus U\notin\mathcal{K}$. The fact that $\phi$ is surjective suggests that there is an $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ such that $\mathcal{K}=\phi(\mathcal{A})$. Thus $L\setminus U\notin\phi(\mathcal{A})$. This means $(L\setminus U)\cap K_{0}=\emptyset$ i.e., $K_{0}\subseteq U$ for some $K_{0}\in\mathcal{A}$. Now suppose that $K_{0}\cap A_{1}=\emptyset$ for some $A_{1}\in\mathcal{K}$. Then $A_{1}\notin\phi(\mathcal{A})=\mathcal{K}$, which is an contradiction obviously. Thus $K_{0}$ intersects each $A\in\mathcal{K}$ and $L$ has the property $(\mathrm{KC})$ as desired.
$(\Leftarrow)$: Obviously, $\phi$ and $\psi$ are order-preserving. Since order-preserving isomorphisms between dcpos are Scott-continuous, the proof will be complete after showing that $\psi\circ\phi$ and $\phi\circ\psi$ are identities, more precisely, the following two equations hold (for arbitrary $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ and $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$):